Status Report

SOHO Daily Meeting Minutes for Friday, 15 February 2002

By SpaceRef Editor
February 15, 2002
Filed under , ,

* For the remainder of the week, call Stein or Emily in case you
need to get hold of a SOC.

* HESSI target: NOAA 9825

Spacecraft Status: SVM Nominal, SUMER OFF
CTOF not yet in nominal operations
Corrective action for thermal reconfiguation
still disabled.

Spacecraft Anomalies: None

Accomplished Activities: VIRGO, CELIAS, NOM HGA Table, Extra VC2

Planned Activities: VIRGO, SWAN, RSL

Upcoming Operations: None

Ground Anomalies:
045/1553 – Wayward command during MDI NRT (8 minute delay)

02/14 DOY 045 – 02/15 DOY 046

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. ICAL, polarity reversal study in the north, Oslo
sunspot study, monitoring studies. Usual weekend studies
will be run.

UVCS: Nominal. Coronal hole observations and synoptic.

LASCO: Nominal. C2/C3 synoptics.

EIT: Nominal. 195 Angstrom CME watch and synoptics, ICAL.

MDI: Nominal. Loads and tests today, plans to get to full disk
magnetograms and dopplergrams campaign today.

TRACE: Nominal. Continue observing AR#9825 and the other regions
within the complex using the UV channel.

SpaceRef staff editor.